OpenStudy (anonymous):

integrate (3sinx-7)cosxdx/4sin^2x+9

OpenStudy (anonymous):

sin x = t ull get (3t-7)dt/ 4t^2 + 9

OpenStudy (anonymous):

now rewrite 3t - 7 = A(8t) + B

OpenStudy (anonymous):

this is integration by parts

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you will get a ln and an inverse tan

OpenStudy (anonymous):

i cant understand.. sorry

OpenStudy (anonymous):

see first substitute

OpenStudy (anonymous):

\[\frac{3t-7}{4t^2 +9} dt \]

OpenStudy (anonymous):

now you split the number up so it looks the derivative of the bottom

OpenStudy (anonymous):

thats why he did 3t-7 = A (8t) +B

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

A= (3/8) B= -7

OpenStudy (anonymous):

\[= \frac{3}{8} \frac{8t}{4t^2 +9} - \frac{7}{4t^2+9} \]

OpenStudy (anonymous):

first fraction integrates to a natural log

OpenStudy (anonymous):

the second to an inverse tangent

OpenStudy (anonymous):

\[= \frac{3}{8} \ln ( 4t^2+9) - \frac{7}{3} \tan^{-1} \frac{2t}{3} +C \]

OpenStudy (anonymous):

fairly sure thats it , then sub back for t

OpenStudy (anonymous):

the arc tangent will look a bit messy when you sub back , but I dont think you can simplify it further

OpenStudy (anonymous):

the arc tan should be (7/6) tan^-1(2t/3)

OpenStudy (anonymous):

poochie, have you done partial fractions?

OpenStudy (anonymous):

its not even partial fractions , you are just spliting up the numerator and getting into a better form